# Toward building next-generation Geocoding systems: a systematic review

### v1

@@ 4.3 Geocoding Interpolation
Linear interpolation. Linear interpolation calculates the output coordinates by linearly scaling the distance between known reference points (e.g., start/end point of a street segmentation). This method assumes that the change between the points follows a straight line, allowing the output coordinates to be estimated proportionally based on the input’s relative position along the line. While this is the most commonly used geocoding interpolation method due to the extensive coverage provided by line-based reference models (e.g., street centerlines), it may yield the least spatially accurate geocodes  [4]. The lower spatial accuracy in linear interpolation stems from several key assumptions, collectively known as the premise of parcel homogeneity  [22]. These assumptions include: (1) the existence of addresses, where parcel counts are solely based on house number ranges; (2) the uniform lot assumption, which presumes that all parcels on a street are of equal size; and (3) the parcel extension assumption, which assumes that parcels completely occupy a street without gaps at corners  [29]. In common linear reference datasets, such as USPS TIGER [^4] https://www2.census.gov/geo/pdfs/maps-data/data/tiger/tgrshp2017/TGRSHP2017_TechDoc_Ch4.pdf, house numbers are often arbitrarily assigned to address ranges for privacy protection, even if they do not exist in reality. Consequently, the linear interpolation algorithm operates on the flawed concept of ”coordinate difference per house number difference” without additional contextual information. Furthermore, geocoding systems often predefine the offset distance from a street centerline to a building without adjusting for local variations, introducing additional errors into the final output  [37].
@@
Area interpolation. Area interpolation uses polygon-based reference datasets to determine a geocoding output point from a given polygon. The main challenge with this method is accurately placing the point to best represent the entire area. Typically, there are three approaches to generating the output point: (1) bounding box, (2) centroid, and (3) weighted centroid  [31]. However, all these methods assume that the derived point accurately represents the building’s location within the polygon. While this assumption holds for small parcels, it often breaks down for larger parcels, especially in rural areas  [7]. Additionally, for irregularly shaped parcels, such as those in ”L” or ”S” shapes, the centroid may not lie within the parcel boundaries  [33].
@@
Point interpolation. Address points, by definition, represent the actual locations of buildings, eliminating the need for interpolation to derive the final geocode  [113]. As a result, this geocoding method is considered to produce the most accurate geocodes. However, reference datasets for address points, which are typically generated either by deriving centroids from building footprints or parcels, or through field collection using Global Positioning System (GPS) technology, are resource-intensive in terms of time and effort  [113, 33]. Thus, such datasets often have limited geographic coverage.
In summary, the spatial accuracy of these three methods is constrained by the available knowledge of building locations and distributions during geocoding interpolations. To address this limitation, one study incorporated auxiliary data , such as online records of building counts and parcel sizes [3], as well as population density  [90]. A later study employed computer vision techniques to detect building rooftops from satellite imagery to determine building distributions and derive centroids for rooftop geocoding. While this approach reduced spatial errors for single-family house parcels, it proved less effective in complex building scenarios.
One future research direction could be an effective interpolation algorithm to leverage multiple data sources, such as building footprints and street view images, to enhance spatial reasoning for geocoding, particularly in areas with multiple buildings or along streets where house number distribution rules are unknown. Additionally, future interpolation methods must resolve geocoding inputs containing spatial expressions, such as distances from a reference point mentioned in Table 1. A recent study  [1] made progress in this area by handling simple distances from POIs. However, further interpolation methods are needed to address more complex descriptions, such as distances related to road intersections or topology associated with highway segments.
5 Opportunities in Geocoding
In this section, building upon the methodologies, challenges, and emerging techniques discussed above, we now explore promising opportunities and directions for advancing next-generation geocoding systems. By synthesizing insights drawn from various geocoding perspectives, we outline pathways for innovation and improvement in future research and development.
@@ 5.1 Uniformed benchmark framework for geocoding
As observed in the previous sections, substantial improvements have been made in the input parsing, storage, and retrieval processes of geocoding workflows. However, a major issue persists: their evaluation results are incomparable due to non-standardized datasets and metrics [52]. In contrast, standardized annotated datasets, like those used in NLP tasks (e.g., NER), are readily available [96]. Additionally, current geocoding evaluation datasets fail to fully reflect real-world input quality, leaving system performance uncertain when handling erroneous inputs. Geocoding inputs often include both syntactic and semantic errors [51], but existing evaluations predominantly use simplistic misspellings and overlook more complex errors, such as semantic inaccuracies.
To address these limitations, we argue that (1) evaluation data should better represent real-world input quality to more accurately assess geocoding performance, and (2) datasets, metrics, and evaluation results should be standardized and made open-access to establish solid baselines for future research. Thus, a potential direction for future work is the creation of such a uniformed benchmark framework that enables ad-hoc evaluation of each sub-process in the geocoding workflow using standardized low-quality input datasets. Recently, [111] introduced a benchmark U.S. address dataset containing over 21 error categories observed in production geocoding services. Further efforts are needed to expand this dataset to have more error types and contain addresses covering a broader range of languages and format variations.
@@ 5.2 A continuously self-improving geocoding system
It is no secret that web search engines can improve the output by (1) continuously crawling web content to maintain the freshness of their reference datasets and provide the latest search results [59], and (2) mining system logs such as input logs, feedback logs to find common input patterns and improve search and recommendation algorithms  [79]. We propose that the next-generation geocoding system should adhere to the principle of continuously updating its reference dataset, enhancing its retrieval capabilities, and retraining machine learning-based modules to ensure high-quality geocoding outputs. This is crucial for the following reasons. First, rapid urban development, such as new construction buildings, often leads to temporal inaccuracies in reference datasets, making them incomplete or outdated. Second, variations in the descriptions or abbreviations of place names, influenced by time and human cognition, can hinder search accuracy; using aliases for these place names has been shown to improve search outcomes, but manually collecting and updating aliases is inefficient. Third, machine learning modules within geocoding systems may experience performance degradation over time, a phenomenon known as model shifting [10]. For example, input errors and variation patterns may deviate from those in the data used to train a machine learning-based address parser, leading to degraded model performance. To mitigate this issue, it is essential to periodically generate a new set of annotated training data and retrain the model to maintain optimal performance. Some works have made some progress in this direction.  [108] applied object detection techniques to geocoding, showing that it can help generate or update address reference data for an input as long as reference records are found for nearby parcels or buildings. [111] developed a framework to analyze input errors and variation patterns from geocoding system logs and generate low-quality geocoding inputs from these patterns.  [16] found that the frequency of erroneous input has a Pareto distribution from logs, and the most frequent spelling variant is directly indexed to reference datasets to improve both the recall and the precision of query results. Future work should focus on designing system architectures and pipelines that facilitate periodic updates of reference datasets, reindexing of specific query fields, and model retraining. Additionally, a robust mechanism to determine the appropriate timing for these updates, along with a corresponding performance evaluation protocol, must be simultaneously developed to ensure the effectiveness of this self-improving mechanism.
5.3 Large Language Models for geocoding
As discussed earlier, LLMs have been applied to create geocoding address parsers, retrieval classifiers, and agents capable of performing geocoding tasks. With the ongoing advancements in LLMs and generative AI, their integration into next-generation systems is expected to bring even greater improvements. We foresee three potential paths for further exploration of LLMs in geocoding. The first path is to use LLM to better comprehend complex and colloquial location descriptions, particularly those involving topological relationships, as outlined in Section 2. These descriptions require parsers not only to recognize address components, but also to extract topological relationships embedded within the descriptions. Meanwhile, such descriptions require geocoding systems to perform more complex spatial interpolations to produce accurate outputs. This means that LLMs must translate the extracted topological relationships into specific spatial calculations. Emerging approaches with LLMs, such as chain-of-thought reasoning [103], present potential solutions to address this challenge, thus defining the second direction for further research. The last direction focuses on leveraging the LLMs’ multilingual and contextual understanding capabilities to overcome the challenges posed by linguistic and cultural diversity in address descriptions. Prior work, such as [105], has demonstrated the effectiveness of using a single neural network architecture with sub-word embeddings to parse addresses across multiple nations. Building on this approach, researchers could explore novel vector representations of addresses, as well as address parsers and retrieval processes that function independently of specific address system rules, potentially paving the way for a universal address search system.
@@ 6 Conclusions
This survey has comprehensively summarized the input and output requirements essential for next-generation geocoding systems, critically examined current geocoding approaches, and highlighted key areas for future advancement. By emphasizing recent breakthroughs in IR, NLP, and LLMs, as well as discussing their promising applications in geocoding, this review seeks to inspire innovative methods and foster interdisciplinary collaboration. Furthermore, it provides a valuable reference for researchers and practitioners aiming to design advanced geocoding algorithms, workflows, and systems, thereby accelerating continued progress and innovation within the geocoding domain.

### v2

@@ 4.3 Geocoding Interpolation
Linear interpolation. Linear interpolation calculates the output coordinates by linearly scaling the distance between known reference points (e.g., start/end point of a street segmentation). This method assumes that the change between the points follows a straight line, allowing the output coordinates to be estimated proportionally based on the input’s relative position along the line. While this is the most commonly used geocoding interpolation method due to the extensive coverage provided by line-based reference models (e.g., street centerlines), it may yield the least spatially accurate geocodes  [boscoe2008science]. The lower spatial accuracy in linear interpolation stems from several key assumptions, collectively known as the premise of parcel homogeneity  [dearwent2001locationalUncertaintyInGeocoding]. These assumptions include: (1) the existence of addresses, where parcel counts are solely based on house number ranges; (2) the uniform lot assumption, which presumes that all parcels on a street are of equal size; and (3) the parcel extension assumption, which assumes that parcels completely occupy a street without gaps at corners  [goldberg2008geocodingbest]. In common linear reference datasets, such as USPS TIGER [^4] https://www2.census.gov/geo/pdfs/maps-data/data/tiger/tgrshp2017/TGRSHP2017_TechDoc_Ch4.pdf, house numbers are often arbitrarily assigned to address ranges for privacy protection, even if they do not exist in reality. Consequently, the linear interpolation algorithm operates on the flawed concept of ”coordinate difference per house number difference” without additional contextual information. Furthermore, geocoding systems often predefine the offset distance from a street centerline to a building without adjusting for local variations, introducing additional errors into the final output  [grubesic2004assessing].
@@
Area interpolation. Area interpolation uses polygon-based reference datasets to determine a geocoding output point from a given polygon. The main challenge with this method is accurately placing the point to best represent the entire area. Typically, there are three approaches to generating the output point: (1) bounding box, (2) centroid, and (3) weighted centroid  [goldberg2013geocodingTechniques]. However, all these methods assume that the derived point accurately represents the building’s location within the polygon. While this assumption holds for small parcels, it often breaks down for larger parcels, especially in rural areas  [Cayo2003]. Additionally, for irregularly shaped parcels, such as those in ”L” or ”S” shapes, the centroid may not lie within the parcel boundaries  [GOLDBERG201239].
@@
Point interpolation. Address points, by definition, represent the actual locations of buildings, eliminating the need for interpolation to derive the final geocode  [zandbergen2008comparison]. As a result, this geocoding method is considered to produce the most accurate geocodes. However, reference datasets for address points, which are typically generated either by deriving centroids from building footprints or parcels, or through field collection using Global Positioning System (GPS) technology, are resource-intensive in terms of time and effort  [zandbergen2008comparison, GOLDBERG201239]. Thus, such datasets often have limited geographic coverage.
In summary, the spatial accuracy of these three methods is mainly constrained by the available knowledge of building locations and distributions during geocoding interpolations. To address this limitation, several studies attempt to incorporate auxiliary data source, such as online records of building counts and parcel sizes [Bakshi:2004:EOS:1032222.1032251], population density  [Schultz_2007], and building footprints detected from satellite imagery  [yin2019DLRooftopGeocoding]. Nonetheless, each of these interpolation approaches exhibits significant limitations when auxiliary data sources are limited or absent.
4.4 Future Directions for Geocoding Components
Based on the analysis of the current state of each geocoding component, we summarize the key challenges and propose paired experimental questions and hypotheses for each component to provoke further investigation.
For Data Storage component, the main challenge is how to make the correct choices to deliver the most effective storage solution that yields high retrieval quality and throughput. To find the answers, we propose the following research questions and hypothesis for investigation.
  • Q1: Which address record segmentation strategy (i.e., address component-based type vs. address component type) should be used for storage?
    H1: Segmenting the address description according to address elements gives more flexibility and leads to better retrieval results.
  • Q2: Which data representation formats (i.e., phonetic encoding, string, dense vector) yield the highest retrieval performance across diverse inputs?
    H2: Dense vector representations outperform string-based and phonetic formats for string-based address elements in general, whereas string-based formats still work well in numeric only address elements such as house number or postal code. especially for noisy or ambiguous inputs.
  • Q3: Which storage media should be used to maximize the system’s retrieval throughput?
    H3: High-performance storage (e.g., in-memory binary file) could potentially increase throughput compared to traditional database techniques.
The first question focuses on the address record segmentation strategy that should be used for storage. Based on the analysis in Section 4.2.1, the segmentation type is highly tied to the parsing results and classification approaches. Thus, we argue that segmenting the address description according to standardized address elements can maximize flexibility since the query build algorithm can combine any of the standardized address elements as needed. The second question addresses the choice of representation formats appropriate for each address segmentation type. Given the fact that geocoding input could have both syntactic and semantic errors, we argue that the choice of representation format should be tailored to each address segmentation. In general, dense vector representations work better on the string-based on address elements such as street- or city-level descriptions, whereas the string representation is more efficient for some numeric only fields by the usage of fuzzy string matching techniques. It is noted that a large portion of this question can be answered by the classifier. The last question concerns how does storage media impact the system’s performance. Given the geocoding reference dataset is considered a fixed and structured data source, if one can compress the dataset and hold them in memory, the query throughput can be significantly increased. All three research questions can be evaluated using the same set of metrics. Specifically, recall and precision can be used to assess retrieval quality, while the number of geocodes processed per second can be used to measure system throughput.
For the Input Parsing component, the main challenge lies in training an address parser that can fully handle the complexity of geocoding inputs. Given the success of modern NLP methods, neural network–based approaches are generally more capable than rule-based or statistical approaches in addressing this complexity. To this end, we propose three research questions to investigate this issue.
  • Q1: Which neural network architecture should be utilized to train an efficient address parser in terms of throughput and parsing quality?
    H1: Transformer-based address parsers offer greater capability than RNN-based neural models but incur higher computational costs during both training and inference.
  • Q2: For the same neural network architecture, does adding non-address short-text context alongside the location description enhance training outcome?
    H2: Embedding the location description with non-address short-text context can lead to improved training performance.
  • Q3: Can a prompt-based address parser backed by generative models outperform the transformer-based address parser?
    H3: Prompt-based address parser requires more efforts to fine-tune a prompt to guide generative models to extract all possible address elements.
The first two questions concern how to train a more robust address parsers. Regarding training data, most existing work relies solely on datasets containing address descriptions, without incorporating additional contextual information. It would be meaningful to examine whether combining fine-grained address description annotations with non-address context could improve training outcomes. [sun2025galloc] developed a platform to annotate location descriptions in disaster-related text messages, which may provide useful initial labeled data. However, training an address parser ultimately requires fine-grained annotations for each address element, as defined in Figure 2. Thus, obtaining such fine-grained training data becomes an essential prerequisite. As for the model architecture, although location descriptions are relatively short, RNN-based models may still experience information loss due to long-sequence dependencies—an issue that can be amplified when the training corpus includes text beyond standard address descriptions. In contrast, transformer-based architectures encode the entire sequence simultaneously and employ multi-head attention to capture intra-sequence relationships more effectively. However, these architectures are also more computationally intensive, leading to more financial costs in the training time and the preparation of the annotated data. These factors underscore the need to evaluate which model is appropriate balanced between model performance and efficiency. Furthermore, incorporating location descriptions from multiple languages would offer valuable insights into how well each model architecture generalizes across linguistic contexts. The last question attempts to evaluate whether or not prompt engineering backed by LLMs can outperform a neural network address parser. We argue that prompt engineering might require more effort to crafted the prompt templates. For example, [hu2023geo] discovered that providing a GPT model with knowledge about location descriptions improves its ability to extract LOC tags and certain standard address elements. Given the fact that more and more foundational LLMs are released, it would be interesting to conduct evaluation comparisons to see if a prompt-based address parser can outperform a neural-network based address parser in terms of parsing quality and the processing latency.
For the retrieval component, a major challenge is to determine which classifier algorithms and representation formats offer the greatest robustness, particularly given the variety of retrieval models currently deployed in production systems, as discussed earlier. Thus, we propose the following research question.
  • Q: Which combinations of classifier algorithms and representation formats provide the most robust retrieval performance across heterogeneous and noisy input scenarios?
    H: Machine learning-based retrieval methods that incorporate learned or hybrid representations (e.g., dense vectors combined with phonetic or string-based features) will achieve higher robustness and retrieval accuracy across diverse input conditions than approaches relying solely on a single representation.
The answer to this question is critical, as it directly impacts the choice of representation formats used to store address records in the reference dataset. We argue that machine learning–based retrieval models generally offer greater robustness because they can adapt to specific geocoding scenarios more effectively than rule-based or probabilistic methods. Additionally, hybrid representation formats can help produce more accurate similarity scores that reflected to the linguistic characteristics of each address element. However, Only through such comprehensive evaluation can one draw reliable conclusions regarding the optimal combination of algorithms and representation formats for next-generation geocoding systems.
The geocoding interpolation component plays a critical role in determining the final geocoded output, as it transforms the matched reference record into geographic coordinates through spatial reasoning. Consequently, the key challenge for this component is to meet the requirements for geocoding outputs, namely, to transparently report the spatial reasoning logic used during interpolation and to quantify the spatial uncertainty associated with the derived location. In this end, a research question can be:
  • Q: How can interpolation methods be designed to enhance their spatial reasoning processes and quantify the spatial uncertainty of the resulting geocode?
    H: Interpolation approaches must integrate more spatial reasoning capability to handle complex spatial expressions, and they must explicitly model proximity to nearby features in order to generate meaningful uncertainty estimates.
We argue that an effective interpolation algorithm must incorporate more advanced spatial reasoning capabilities to handle the complexity of geocoding scenarios such as areas with unknown building distribution patterns or location descriptions containing topological relationships. Recent work by  [al2019towardsGeocodingSpatialExpressions] made some progress in this direction by addressing simple distance-based expressions relative to points of interest. Another important aspect that new interpolation methods should have is to trace and document their spatial reasoning steps so that the interpolation procedure can be transparently reported and the uncertainty of the final output can be appropriately estimated.
In this part, we present a set of research questions and hypotheses for each geocoding component. We envision that this list will help motivate future experimental investigations and guide the development of next-generation geocoding techniques.
5 Pathways Toward Next-Generation Geocoding Development
Building on the methodologies, challenges, and remaining research questions discussed above, we now explore broader opportunities for advancing next-generation geocoding systems. In contrast to the component-level research directions outlined earlier, the opportunities presented here highlight new pathways or new paradigm for future geocoding system development.
@@ 5.1 Uniformed benchmark framework for geocoding
As discussed in the previous sections, developing new geocoding techniques or identifying optimal solutions requires systematic performance evaluation. However, a major issue persists: existing evaluation results are often incomparable due to the absence of standardized datasets and metrics [jacquez2012research]. By contrast, many NLP tasks, such as NER, benefit from well-established and standardized annotated datasets [conll2003NERdataset]. To date, existing geocoding evaluation datasets fail to fully reflect real-world input quality. Although geocoding inputs often include both syntactic and semantic errors, existing evaluations predominantly use simplistic misspellings and overlook more complex errors, such as semantic inaccuracies, leaving system performance uncertain when handling erroneous inputs. Recently, [yin2023chatgpt] introduced a benchmark U.S. address dataset containing over 21 error categories observed in production geocoding services. However, further efforts are still needed to expand this dataset to encompass a wider range of error types, languages, and format variations.
To benefit future development of geocoding systems, we argue that the foremost priority is to create a unified benchmark framework for geocoding. This framework should have the following features: (1) evaluation data should better represent real-world input quality to more accurately assess geocoding performance; (2) datasets, metrics, and evaluation outputs should be standardized for both component-level and end-to-end geocoding workflows; and (3) the framework should be made openly accessible to permit future researchers to evaluate individual geocoding sub-workflow or full geocoding systems as needed.
@@ 5.2 A continuously self-improving geocoding system
It is no secret that web search engines can improve the output by (1) continuously crawling web content to maintain the freshness of their reference datasets and provide the latest search results [lewandowski2008three], and (2) mining system logs such as input logs, feedback logs to find common input patterns and improve search and recommendation algorithms  [pan2015feedbackUpateRecSystem]. We propose that the next-generation geocoding system should adhere to the similar principle to continuously update its reference dataset, enhance its retrieval capabilities, and re-train its machine learning-based modules to ensure high-quality geocoding outputs. We named this geocoding system as a self-improving system. This is crucial for the following reasons. First, rapid urban development, such as new construction buildings, often leads to temporal inaccuracies in reference datasets, making them incomplete or outdated. Second, variations in the descriptions or abbreviations of place names, influenced by time and human cognition, can hinder search accuracy; using aliases for these place names has been shown to improve search outcomes, but manually collecting and updating aliases is inefficient. Third, machine learning modules within geocoding systems may experience performance degradation over time, a phenomenon known as model shifting [chen2021modelshifting]. For example, new input errors and variation patterns may deviate from those in the data used to train a machine learning-based address parser, leading to degraded model performance. Several works have made some progress to mitigate these issues. For example,  [yin2019DLRooftopGeocoding] applied object detection techniques to geocoding, showing that it can help generate or update address reference data for an input as long as reference records are found for nearby parcels or buildings. [yin2023chatgpt] developed a framework to mine input errors and variation patterns from geocoding system logs and generate low-quality geocoding inputs from these patterns.  [clemens2018indexingSpellErrorforGeocoding] found that the frequency of erroneous input has a Pareto distribution from logs, and re-index the most frequent spelling variant into reference datasets to improve both the recall and the precision of query results. We argue that by leveraging historical geocoding transactions to automatically re-train an address parser and a classifier or to augment the reference dataset, it is possible to achieve a self-improving geocoding system.
However, to fully deliver such a geocoding system, several questions must be addressed in advance. First, when mining geocoding logs, sensitive information, such as contact details or personally identifiable attributes, must be carefully removed to prevent potential privacy leakage. Therefore, a mechanism to perform for K-anonymity or differential privacy needs to be embedded into the workflow so that sensitive information is neither used during model training nor re-indexed into the reference dataset. Second, we need to have a methodology about how to update each geocoding component in a systematical manner to improve overall outcome. For example, we need to have clear answers for the following questions: (1) to what extent can retraining the address parser enhance parsing outcomes, and (2) are there more cost-effective alternatives, such as reorganizing the reference dataset that could achieve similar improvements? Third, once the update strategies are defined, the next challenge is determining when these updates should occur. Given the substantial computational and temporal costs of model training, it is critical to develop a performance evaluation protocol that triggers the self-improvement process only when necessary, thereby ensuring the improvement of the whole geocoding system is achieved at minimal cost.
5.3 LLMs for an end-to-end geocoding workflow
In previous section, we observe that LLMs have been applied to develop address parsers, retrieval classifiers, and agents capable of performing geocoding tasks. Moving beyond these improvements to individual components, we envision that LLMs will increasingly support geocoding tasks in a more end-to-end manner. Specifically, for the geocoding textual retrieval phase, LLMs offer an opportunity to shift from traditional address element-based matching to an embedding-driven retrieval framework. Instead of segmenting the input description into indexed address elements and searching the reference dataset accordingly, LLMs can generate a vector representation of the entire address description and retrieve the most similar record directly from the reference dataset. Furthermore, textual similarity and spatial distance can be jointly incorporated into a loss function to learn embeddings that align both linguistic and spatial proximity . Under such a framework, when a new address description is received, the system would encode the full description into an embedding and then identify the most similar address record based on these learned representations. For the geocoding interpolation workflow, we expect that LLMs can help translate topological relationships into specific spatial calculations, especially, using chain-of-thought reasoning [wei2022chain] to derive a set of step-by-step spatial reasoning procedures. To this end, we can have an end-to-end geocoding workflow that can minimize error accumulation across components and improve adaptability to diverse geocoding inputs. Needless to say, pursuing this direction requires substantial effort in model training, validation, and continuous evaluation. For practitioners, it is therefore essential to carefully assess the trade-offs between the computational and operational costs and the potential performance gains of adopting such an approach. However, this remains an interesting direction that warrants further investigation by future researchers.
@@ 6 Conclusions
This survey has outlined the fundamental requirements that modern geocoding systems must satisfy and assessed how well current approaches address those needs. By examining the major functional elements of the geocoding workflow, we highlighted the methodological foundations of each component and identified the gaps that remain. We then formulated targeted research questions and hypotheses to guide systematic evaluation and to support the development of more capable geocoding solutions. Building on these analyses, we also discussed broader opportunities and future directions for advancing the field.
Looking ahead, we anticipate that the evolution of next-generation geocoding systems should follow two major trajectories: enhanced understanding of location descriptions and improved spatial reasoning. The first trajectory should involve incorporating semantic representations, neural parsing models, and hybrid similarity measures to robustly interpret diverse and complex address expressions across languages, formats, and domains. A related extension of this trajectory is to learn joint embeddings for address records that reflect both their textual descriptions and their spatial and topological relationships, enabling retrieval models to capture linguistic similarity and geographic proximity in a unified representation space. The second trajectory emphasizes deeper spatial reasoning by integrating geometric context, boundary awareness, and uncertainty estimation to derive multi-step spatial reasoning procedures that more faithfully reflect the distribution of objects. We invite the research community to pursue these trajectories to advance geocoding systems that not only yield more accurate spatial coordinates, but also better understand and interpret real-world spatial complexity.